OpenStudy (anonymous):

How would you solve this system of equations: 0.5x-4=y and -2.5x-2.5=y ?

OpenStudy (anonymous):

Do you need to do it by graphing?

OpenStudy (anonymous):

Yes. I'm getting confused by finding these equations slope & y intercept, and because of the decimals.

OpenStudy (anonymous):

On these it might be better to do the x and y intercept method

OpenStudy (anonymous):

Oh. Well, The x on the first one would be 0.5 and the x on the second one would be 2.5, I think?

OpenStudy (anonymous):

To get the x intercept substitute 0 in for y and solve for x 0.5x-4=y .5x-4=0 .5x=4 x=8 so the x intercept for that one would be (8,0)

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I put 0 where y is and solved for x

OpenStudy (anonymous):

Thats how you get the x intercept

OpenStudy (anonymous):

I'm confused. I know it works out, 0.5*8-4=0 , but how do you work that out to get 8?

OpenStudy (anonymous):

Alright we have 0.5x-4=y put 0 where y is .5x-4=0 add 4 to both sides .5x=4 divide both sides by .5 x=8

OpenStudy (anonymous):

And thats the x intercept so (8,0)

OpenStudy (anonymous):

Ok, I get it. So, on the second equation, the x intercept would be (5,0)

OpenStudy (anonymous):

Not quite, but lets finish the first one first i dont wanna get mixed up lol. To get the y intercept you do the same thing except instead of putting 0 in for y, you put it in for x. 0.5x-4=y .5(0)-4=y -4=y so y intercept would be (0,-4) now you can graph that line using the x and y intercept

OpenStudy (anonymous):

Haha. :) Ok, I get the y intercept part. Thats pretty simple. But the x intercept is a lil more confusing. I think the y intercept on the second equation would be -2.5 because -2.5(0)-2.5= -2.5 but I don't get the x intercept part

OpenStudy (anonymous):

Yep, so you have one line plotted with the x and y intercept (0,-4) and (8,0) You're right on the next ones y intercept so (0,-2.5) x intercept would be -2.5x-2.5=y -2.5x-2.5=0 what do you get when you solve that for x?

OpenStudy (anonymous):

I don't know. How do you solve for it?

OpenStudy (anonymous):

Its a simple equation, first add 2.5 to both sides to get -2.5x=2.5 then divide both sides by -2.5 x=-1

OpenStudy (anonymous):

Oh, I understand. Thank you! Your a genius. :)

OpenStudy (anonymous):

It would be (-1,0) right?

OpenStudy (anonymous):

Yep, now find where they intersect to get your solution. You're welcome